7.2
Use the following composition of expenditure for the economy of a country named the Republic of Tapuwa, for the year ended 2019 to answer the questions below:
COMPONENT
R millions
Consumption expenditure (C) Investment (I)
Government spending (G) Exports (X)
Imports (Z)
Depreciation
Foreign payment to the rest of the world
Foreign payment from the rest of the world
Show all calculations and formulae.
9 000 6 500 7 000 1 800 2 400 700 300
250
© The Independent Institute of Education (Pty) Ltd 2020
Page 9 of 10

20
2020
Q.7.2.2 Q.7.2.3 Q.7.2.4
Compute the value for the country’s GDP (Gross Domestic Product) (2) at market price.
Determine the value of the country’s NNI (Net National Income) at (3) market price.
If it is predicted that the GDP will increase to 22 000 in 2020, (3) calculate the growth rate between 2019 and 2020.
END OF P
1
Expert's answer
2020-11-30T16:29:50-0500

a) ANSWER


GDP = $21,900


SOLUTION

GDP at market prices

= C + I + G + Net Exports

= C + I + G + (X - M)

= 9,000 + 6,500 + 7,000 + (1,800 -2,400)

= 22,500 + (-600)

= $21,900



b) ANSWER


NNI = $21,150


SOLUTION

NNI at market prices

= GNP - depreciation

= GDP + Net property factor payments from abroad - depreciation (capital consumption)

= GDP + (Foreign income - Foreign payments) - depreciation

= $21,900 + (250 - 300) - 700

= $21,900 + (-$50) - $700

= $21,150



c) ANSWER


Growth rate = 0.457%



SOLUTION

"Growth \space Rate = \dfrac { 2020 \space GDP - 2019 \space GDP}{2019 \space GDP}× 100\%"


"Growth Rate = \dfrac {22,000-21,900}{21,900}×100\%"


"Growth \space Rate = \dfrac {100}{21,900}×100\%"


"= 0.00456621×100\%"


"= 0.466621\%"

"= \bold {0.457\%}"